Class Discussions on BibLit

Comp Lit 280 December 1996

Chapter 1

Texts and Contexts - Week 1

Augustine, On Christian Doctrine

Acts 2:1-36; 7:1 to 11:30; 13:17-52; Romans 9:1 to 11:36; I Cor. 1:1 to 4:21; 10:1-13

1.1 Augustine Summa Theologica

Since there was little discussion on this text, I have included Augustine’s comments on interpretation. For a brief discussion, see Section 6.9 on page 25.

Summa Theologica I, 1, 10^{1}

Whether in Holy Scripture a word may have several senses?

Objection 1. It seems that in Holy Writ a word cannot have several senses, historical or literal, allegorical, tropological or moral, and anagogical. For many different senses in one text produce confusion and deception and destroy all force of argument. Hence no argument, but only fallacies, can be deduced from a multiplicity of propositions. But Holy Writ ought to be able to state the truth without any fallacy. Therefore in it there cannot be several senses to a word.

Objection 2. Further, Augustine says (De util. cred. iii) that “the Old Testament has a fourfold division as to history, etiology, analogy and allegory.” Now these four seem altogether different from the four divisions mentioned in the first objection. Therefore it does not seem fitting to explain the same word of Holy Writ according to the four different senses mentioned above.

0This document includes Miceal Vaughan’s and Gabrielle Shearer’s responses to questions presented on the BibLit Discussion groupe. Prepared with L^{A}TEX

1Hypertext Version Copyright ) 1995 Kevin Knight

Objection 3. Further, besides these senses, there is the parabolical, which is not one of these four.

On the contrary, Gregory says (Moral. xx, 1): “Holy Writ by the manner of its speech transcends every science, because in one and the same sentence, while it describes a fact, it reveals a mystery.”

I answer that, The author of Holy Writ is God, in whose power it is to signify His meaning, not by words only (as man also can do), but also by things themselves. So, whereas in every other science things are signified by words, this science has the property, that the things signified by the words have themselves also a signification. Therefore that first signification whereby words signify things belongs to the first sense, the historical or literal. That signification whereby things signified by words have themselves also a signification is called the spiritual sense, which is based on the literal, and presupposes it. Now this spiritual sense has a threefold division. For as the Apostle says (Heb. 10:1) the Old Law is a figure of the New Law, and Dionysius says (Coel. Hier. i) “the New Law itself is a figure of future glory.” Again, in the New Law, whatever our Head has done is a type of what we ought to do. Therefore, so far as the things of the Old Law signify the things of the New Law, there is the allegorical sense; so far as the things done in Christ, or so far as the things which signify Christ, are types of what we ought to do, there is the moral sense. But so far as they signify what relates to eternal glory, there is the anagogical sense. Since the literal sense is that which the author intends, and since the author of Holy Writ is God, Who by one act comprehends all things by His intellect, it is not unfitting, as Augustine says (Confess. xii), if, even according to the literal sense, one word in Holy Writ should have several senses.

Reply to Objection 1. The multiplicity of these senses does not produce equivocation or any other kind of multiplicity, seeing that these senses are not multiplied because one word signifies several things, but because the things signified by the words can be themselves types of other things. Thus in Holy Writ no confusion results, for all the senses are founded on one—the literal—from which alone can any argument be drawn, and not from those intended in allegory, as Augustine says (Epis. 48). Nevertheless, nothing of Holy Scripture perishes on account of this, since nothing necessary to faith is contained under the spiritual sense which is not elsewhere put forward by the Scripture in its literal sense.

Reply to Objection 2. These three—history, etiology, analogy—are grouped under the literal sense. For it is called history, as Augustine expounds (Epis. 48), whenever anything is simply related; it is called etiology when its cause is assigned, as when Our Lord gave the reason why Moses allowed the putting away of wives—namely, on account of the hardness of men’s hearts; it is called analogy whenever the truth of one text of Scripture is shown not to contradict the truth of another. Of these four, allegory alone stands for the three spiritual senses. Thus Hugh of St. Victor (Sacram. iv, 4 Prolog.) includes the anagogical under the allegorical sense, laying down three senses only—the historical, the allegorical, and the tropological.

Reply to Objection 3. The parabolical sense is contained in the literal, for by words things are signified properly and figuratively. Nor is the figure itself, but that which is figured, the literal sense. When Scripture speaks of God’s arm, the literal sense is not that God has such a member, but only what is signified by this member, namely operative power. Hence it is plain that nothing false can ever underlie the literal sense of Holy Writ.

Ordered Cosmos - Week 2

and Disordered Humanity

Boethius, Consolation of Philosophy; Dante, Inferno Cantos 1-5; Paradiso 28-29

2.1 Boethius

I have no quarrel with your claim that “Boethius does conform to Christian doctrine,” and I hope I never suggested otherwise. Your questions raise a number of fundamental issues, ones which others have raised in class, so I am posting them along with my responses to our discussion list. You articulate your views clearly and with real sophistication (and force), so I am confident you can deal with my responses in kind.

My specific comments are provided below, following parts of your text. I’m not sure I read the last sentence as demanding the existence of “some

OTHER incomprehensible being.” What pushes you to that view? I read the “judge” as the same God Boethius has talked about earlier, who is by no means “incomprehensible.” In fact, I would say that Boethius insists that his god is fully comprehensible, even by a character like the depressed first-person speaker of this work.^{1}

1Sarah Fisher wrote: (Either you or the introduction, I’m not sure I remember, stated that The Consolation of Philosophy was written early on during his imprisonment; Boethius probably had not yet comtemplated theological aspects fully.) Boethius may have been a Christian, but the Consolation is fully dependent upon the rationale that all things must be proven in order to be believed; this rationale contradicts one of the main tenets of Christianity: faith–which I think is belief without tangible proof. I am content with my own personal opinion about wisdom vs. philosophy

There seems to be little doubt that Boethius wrote a number of theological tracts (including one on the Trinity), and not all his writing was done in prison (if, in fact, ANY was), so I have real difficulty with your version of his intellectual and faith history. If he DID write the Consolation in prison, awaiting his execution, then this is AFTER he has written a number of orthodox theological tracts about Christian beliefs.

Does that throw an obstacle in the path of your “personal opinion” in this case? I don’t accept, by the way, your above definition of “faith” as a tenet of Christianity. It may be the definition accepted in YOUR Christianity, but it will not stand as an accurate (or stable) definition for all Christianity, in all its varied historical and denominational versions.

The views^{2} you express here are consistent with a range of modern (somewhat anti-rationalist) views of the relations between philosophy and theology, reason and faith. But the point I’ve been trying to make is that this modern (religious) distinction and the valuation of the terms (which is the other side of the rationalist’s view that religion is superstition) is not at all shared by the late-antique and early medieval writers we are currently reading. Nor will that distinction become dominant at any point during the Middle Ages. It is not until the very late Middle Ages that it starts to rise as a widespread belief and it must wait for the Reformation to give it rich soil on which to grow.

If you have grappled with Augustine’s On Christian Doctrine^{3}, you will quickly have recognized the rich conflation of “philosophy” and religion.

Also, while you are correct that monks separate themselves from the world, it is NOT primarily “worldly” reason/philosophy they are running from. In fact, monastic centers are where the educational ideals and practices of Graeco-Roman schools (including reason and philosophy and their “pagan” texts were preserved). You are correct that monasteries did provide an environment for mystics and others who sought what you characterize as “wisdom,” but it was not everyone who did that, nor was it fundamental to the monastic way of life (as you will see, I trust, in the Rule of St. Benedict^{4}).

2Ibid: Philosophy involves study or contemplation driven by the humanly selfish desire to KNOW all things and their purposes through Reason; philosophy strives for an ultimate answer, derived logically within the comprehensive ability of man. Yet Wisdom, as shown in being personified in the Wisdom books of the Bible and as being the Word which was and is God, is the characteristic of a man who understands that he cannot understand all things, considering that the intellect of God is as imcomprehensible as infinity. Wisdom becomes the goal of the monks, who eschew worldly ways and philosophies (!), in the attempt to understand as much of God as is possible; monks follow God because they know God is omnipotent and omniscient, claims which no human could ever make. Therefore, I see philosophy as the inane attempt to create a logic which explains all things and wisdom as the sagacity to realize humans cannot understand everything while in their present capacity on earth.

3see section on Ordering Existence

4Section 3.1 on Page 7.

Law Divine and Human - Week 3

Reordering Human Disorder. The distinction between secular and religious authority.

3.1 Rule of Saint Benedict

Rules for Children

From what I know, this^{1} practive of people offering their children to monasteries was a common (and perhaps increasingly common) practice. For one thing, monasteries were about the only places where anything like “education” took place in the early Middle Ages: they had libraries and were committed to transmitting the knowledge contained in them (and not always only of the obviously religious texts). It was also the case that such training could lead to careers and promotion in the Church–or even service in secular coursts–which was not a neglible matter in terms of career planning.

On the other hand, having a relative in the monastery to pray for you and your family had its attractions. And finally, monasteries were for the most part secure and civilized places, which often became so “comfortable” that they inspired repeated reforms to try and call them back to their “religious” and “spritual” functions.

1On Tue, 15 Oct 1996, M. Bowers wrote: In “The Rule of St. Benedict” St. Benedict lays out the rules for the monks in the monastery. I was surprised to read of the rules for children in the monastery. Towards the end of the book he talks about noblemen or poor offering their children to the service of God. How common was this practice and were children placed in the monasteries because their parents were not able or did not want to care for them?

All in all, then, there might be any number of practical and religious reasons why noblemen (and others) might seek to have children of theirs admitted to monasteries.

3.1.1 Obedience

I’d certainly take the coexistence of the two interpretations^{2} as the most likely. Most hierarchical systems (including most of the medieval versions) tend to authorize the power of leaders by associating them with God (or the gods). One hears about the “divine right of kings” and Gabrielle’s remarks earlier in the week about the “priestly,” even divine, aspect of Byzantine emperors are worth remembering.

Most “obedience” is related to some differences in power between the two parties (the obeying and the obeyed), and at least some of that power can be abstract or conceptual (rather than merely physical). We talk of people today as having “charisma,” i.e., some undefinable power that causes others to want to do what s/he says. That word is directly related to the concept of “grace” (that is, some spiritual or divine force to which a human individual has access). That can come by virtue of one’s position–or it can be the result of some innate qualities of the person.

Clearly in the monastic setting, the position of the abbot is what accords him the grace/power which is identifiable with divine authority. HOW one got to BE an abbot may depend, of course, on more innate qualities of person. For the most part, however, the obedience is owed to what is right and proper, and it is an assumption that God is to be identified with that, and that the more immediate “rulers” on earth (parents, lords, kings, abbots) derive their authority from those sources. Indeed, one of the contributions of later medieval thinkers (like Marsilius of Padua) is to define the grounds on which a figure of earthly authority could be deposed, and to evolve theories of governance which depend on the will of the people as defining what is right and proper for their community.

A long and rambling answer, I know. But this is another case where OUR modern, American assumptions/views are radically different from those of earlier European cultures like the ones we are studying.

2R. Thompson: “I’m a little confused on how to interpret the word ’obedience’ when read in The Rule of St. Benedict. Is it refering to obedience to God, or obedience to God through the monasterie’s abbot? Or is the meaning a combination of the two interpretations?

3.1.2 Monasteries and Economies - From G. Shearer

I wanted to mention that monasteries were also significant landholders and could provide a fairly significant economic base for people in the surrounding region. If you think of a monastery as a local big business as well as a religious order, you get an idea that it was a viable part of the community. In other words, monasteries were more high profile in the Middle Ages. I was thinking of this during Miceal’s lecture yesterday–we would probably think it strange if one of our relatives moved to Lacey and joined the Benedictine Monestary there, but that is because monastic orders are anachronisms today. However, we do have a number ofbusinesses in our area that suggest economic stability, security, social status, community, etc. Imagine MicroSoft with a theological bent, and you get the idea. Abbot Bill Gates. Any others have an opinion re: monasteries? Ive noticed from your Monday questions that quite a few of you are interested in this, and this might make a nice way of discussing it. I’d love to see this listproc account used for something besides filling the five question quota. We dont have to be formal, and its a good way to try out ideas/share info, etc.

By the way, if you come across anything good on the Web pertaining our interests, let us know. There is a lot of weird and wonderful material out there.

Journey(s) of Life - Week 4

Quest of the Holy Grail; “The Wanderer”; “The Seafarer”

Genesis 37:2 to 50:26; Exodus 1:1 to 18:27

4.1 Relics and Quest of the Holy Grail

We’ve not really spent THAT much time on relics, but in the context of this week’s discussions of pilgrimages we will be taking it up. Since others may have these questions (specific or general), I will forward this reply to the discussion list.

There are really a number of different sorts of things that are collectively called “relics”–and they are sometimes catergorized into distinct “classes” (first, second, etc.).

• A “first-class” relic would be one which is actually a part of a saint’s body (the entombed remains, for example, or a single piece of the body: a bone, a head, an arm): these were particularly venerated as “holy” in the Middle Ages, and were often the object of pilgrimage.

• A “second-class relic” would be some object (a piece of clothing, say, or a tool) owned by and used by a saint.

• A “third-class relic” (probably a much later development, when “relics” had become really popular objects of veneration) would be something (a piece of cloth, etc.) which had touched a first-class (or secondclass) relic directly–in other words, something that had traveled to a site where some particularly important relic existed and had come back with some of its “power.”

In the “Quest,” the relics you mention would probably fit in one or other of these categories of relics (most likely first or second). The belief that such relics were holy and worthy of veneration is very like what Gabrielle had to say about icons. And, indeed, the attitudes toward relics also have the strong division of opinion that we can see during the so-called Iconoclastic controversies in Byzantium: some firmly believe in them, while others think they are obstacles to true religion. ^{1}

1Mon, 21 Oct 1996, Courtney Dawn Cascio wrote: “First instance is when Galahad and the other knights knelt and swore before the relics. The other time is when the ’white knight’ is telling the story of the shield on page 59 he states that Evalach asked Josephus for a relic of his life. So were the relics that the knights swore before possessions of others that had been deemed ’holy’? Or are they in no way related, but two definitions for the same word?”

Modeling Kingship - Week 5

Einhard’s Life of Charlemagne; Song of Roland

Genesis 12:1 to 25:18; Judges 2:6 to 5:31; 13:1 to 16:31(Samson); I Sam 8:1 to II Sam 2:4 (Saul and David); II Sam 11:1 to 19:9; I Kings 1:1 to 5:14; 9:26 to 13:34

5.1 Life of Charlemange

5.1.1 Order and the Individual

Not easy questions ^{1} to answer–as you probably know. My take on it is that there wasn’t as much “information” available in the Middle Ages, that is, not much in books and other repositories (besides the individual’s experience and memory). Since it wasn’t available, no one really felt left out of it. The very few who could read and have access to books had some power over others, but it may not have counted much against physical power and bodily health.

In the high middle ages there was an information explosion of sorts and increase access to education (universities, etc.), and the educated classes got bigger–and were not always fully utilized in the church or state government. To an extent, the Renaissance and Reformation are the product of the increasing education and urbanization of European society, and efforts by the educated to locate new places for “work” lead them to deal with other parts of society and the world than might have attracted efforts earlier. This probably accounts, at least in part, for the rise of “science” and other enterprises–and their effects on social practices and understanding.

1On Mon, 28 Oct 1996, R. Hammond wrote: 1)Latter in the Middle Ages, do people grab a greater sence of order and understanding? It seems there is a lot of pure intent by idividuals of society, yet their effect is limited due to an enormous amount of limited knowledge. Or is it this lack of knowledge that leads to the Reformation, or the Enlightenment. Do you think this struggle to understand in the past has brought our world culture to the “prove it” level of acceptance?

I’m not sure how to measure “changes in ... mentality. ^{2}” There certainly were cultural and practical changes in people’s lives. Do you see or suspect some particular differences in attitude between Boethius/ Benedict and Einhard/Notker?

Yes, I wouldn’t want to oversimplify and homogenize something as large and variable as human culture. There were no doubt differences ^{3} , but the question is really where did the differences show up? There were, I think, large areas of philosophical and practical agreements that would allow us to talk somewhat about the Middle Ages as if they were some single era distinct in important ways from the “modern” and the “ancient.” The exact nature of those differences might only obscure more subtle, but nonetheless important, differences WITHIN the broad historical and geographic bounds of the European Middle Ages.

A branch of the Western Germanic invaders of Europe. They^{4} conquered and ruled the area that is now northern France and NW Germany, but extended their control over much of the surrounding area–as far as N Italy and Austria and S and E France and Belgium.

The sacrament of matrimony ^{5} was a very late “invention.” The Church was not as involved in regulating marriages in Charles’s time, and it tended to be a legal contract (which could be broken for cause). If I remember correctly, it is not until about the 12th century that the number seven is settled on for the number of sacraments, and matrimony only makes it into the list at about the same time.

2Ibid. 2)Have there been any major changes in the mantality of people from the fifth and sixth century, that are important in understanding The Lives of Charlemagne?

3Ibid. 3)Is the mind set of cultures in the Middle Ages similar all through out Europe?

I realize there had to of been some differences, but are there differences between Areas like Modern England and France that can facilitate the significance of the The Quest and Charlemagne?

4Ibid. 4)Who where the Franks?

5Ibid. 5)When if says that Charlemagne “dismissed” his first wife, what does that mean? I always understood that marriage at at that time was not a terminating thing.

Divorce, I understood, was not an option.

5.1.2 The Role of Bishops and Priests

Bishops and priests may, in fact, also have been warriors ^{6} earlier, but certainly by the time of the crusades (and the Song of Roland is almost exactly contemporary with the First Crusade) clergy probably did indeed fight. It was, after all, a holy war and so the division between secular and religious interests were not easily separated. The Roland clearly imagines this as a war between good and evil, so having an Archbishop killing evil men poses no moral problems.

And lest you simply see this as “poetic license,” you should know that quasi-monastic orders, like the Knights Templar, were formed explicitly of knights to protect holy sites and transportation routes during the period of the crusades.

Your^{7}, questions raise fundamental issues, ones which go to the very heart of issues I have been trying to get across during these first few weeks. Many of the strongly held beliefs of post-Reformation Christians are radically different from those professed by pre-Reformation, medieval Christians. There are significant differences in assumptions and emphasis, and they can only be understood when each is given the opportunity to be heard fully. After that stage, the arguments can begin in earnest.

He (Dante)got it from late medieval theologians like Thomas Aquinas and others.

I’m not sure “rationalize^{8}” is the term to use; he is attempting to describe in imaginative, fairly concrete form what the ideas he and contemporaries held to be true.

Medieval Christianity (and even preChristian Judaism–see II Maccabees 12:43-46) believes that actions/prayers in this life can affect the conditions of those already dead. The story of the rich man and Lazarus seems to draw on such ideas, for example. Anyway, if there is improvement possible after death (and commentary on Matt 12: 32 by Augustine and others makes this explicit), then there must be some middle state between Hell and Heaven, one which is not likely to be eternal, as the other two are.

6On Mon, 28 Oct 1996, J. Adams wrote: “Was it considered appropriate for an archbishop to go into battle - and in battle, to kill?”

7On Mon, 28 Oct 1996, J. Leggett wrote: “I am not quite sure where Dante’s whole idea of Purgatory came from.”

8Ibid. How does Dante rationalize all of His levels and where does he get the idea of ’earning’ your way into heaven when the Bible says in Romans chapter 10 verse 9 ’That if you confess with your mouth, ’Jesus is Lord’ and believe in your heart that God raised Him from the dead, you will be saved.’ then following in verse ten it says ’For it is with your heart that you believe and are justified, and it is with your mouth that you confess and are saved.’

Catholic theology makes distinctions among sins (mortal vs. venial) and pretty much holds that everyone sins in some way or other and most die with some stain of sin on their soul. Is it just, the theologians ask, for God to send all of these to hell, when their sins are not all equally grievous? What do you do with a venial sinner, who’s not ready for Heaven but not deserving of Hell.

Out of thinking and commentary of this sort evolved the belief in a middle place where sinners can become purified of their residuum of sin. It’s not a matter of “earning” Heaven, then, but of cleansing one’s self to be worthy of admission that lies at the basis of the theology of Purgatory.

The Middle Ages (and modern Roman Catholicism) do not put as much stress on the concept of “justification” that you point to. And when they do, they insist that one’s faith, belief, and confession must necessary express itself in one’s behavior and actions: the fullest meaning of one’s words are found in one’s life. And while we can state high principles and ideals, we can seldom live every moment in full accord with them. That limitation and repeated failure should not, however, be a source of despair (or final condemnation) and a truly just God would also include a component of mercy that allows the fallen to rise again.

5.2 Song of Roland

The appointment of bishops has gone through a number of different formats over the years. Originally, apparently, they were basically elected by the members of the community and then installed by the laying on of hands by another bishop. Presently, in the Roman Church, bishops are nominated and appointed by the central church organization in the Vatican and officially named by the Pope. That is, however, a practice that has had its ups and downs even in fairly recent times, and it certainly was not a common arrangement in the Middle Ages–especially not in established Christian areas. There, the local powers-that-be would play an important role in nominating and even in appointing bishops, sometimes without (or even against) the wishes of the Bishop of Rome. Some strong Popes could assert their primacy in these matters, but so could some strong kings, dukes, and lesser lords.

The Pope did, of course, appoint bishops and usually was the prime authority in cases of “missionary” areas, like early England or Merovingian France. But once the local secular authorities became Christian, they more likely played a big role in subsequent appointments. Beginning in the eleventh century or so, there were major disagreements between secular and ecclesiastical authorities over matters like this and a division of church and state authority in the appointment of bishops was a major element of those debates and of the reforms they brought about. IN the later Middle Ages, then, secular rulers played a less prominent and absolute rule in appointing bishops–though, of course, they would still have an important voice (as they sometimes still do) in the choice of important ecclesiastical officials in their territories.

Your question^{9} about Saxon devil-worshippers points to a tendency in medieval writings dealing with crusades (or their ilk, such as this imperial expansionism of Charlemagne) to demonize the enemy. This is, of course, not limited to medieval Christian writings: most wars tend to locate “us” on the side of the angels, and the enemy on the side of the devils. The matter of “conversion” is of course a problematic one, particularly since we take the word’s meaning primarily to refer to things like “belief” and “feeling.” We should, I suppose, broaden those terms to include “institutional affiliation” as an important factor. This sort of political affiliation is probably a more significant feature of medieval conversions than would be the case today, perhaps. If the king or ruler of a tribe “converted” to Christianity, it was assumed that all his “people” did also. It is unlikely, however, that even when the king converted, he had gone through some profound intellectual or spiritual “conversion” and instead made a commitment to belong to the church, perhaps accepted baptism as a sign of that commitment, and then would begin to adjust his life and his beliefs as he became educated in what it meant to be a Christian. The symbols of belonging, and the political costs/benefits, would be more important measures of conversion than some interior change of heart.

In such a context, I think your question takes on a rather different coloration. Does it change your thinking about the question of “force” in such a case as that of the Saxons, and affect your sense of what WOULD be the measures of a “genuine” conversion in such circumstances?

9On Mon, 28 Oct 1996, J. Sullivan wrote: “How is it that ‘devout clergy’ would encourage Charlemange to put aside his wife ‘As though she were already dead?’ How is it that Charlemange can appoint bishops in the Roman church? Why does Einhard keep refering to the Saxons as devil worshippers and what effect on the sincerity of their Christianity might the fact that they were forced to become Christians have?”

Love Songs - Week 6

The Sexual, The Sacred, and The Courtly Dante, Vita Nuova; Chaucer, Miller’s Tale Proverbs 3:13ff; 8:1 to 9:18; Song of Songs; Hosea (and cf. Ezech. 16:1-

63)

6.1 Dante

I don’t know that all (or even many) Dante scholars would accept this^{1} view. It is, however, a possible “source”–even if the direct connections cannot be made between the Islamic accounts and Dante’s. I don’t know the former, but my understanding from what I’ve read about this theory is that the similarities are of only a fairly general sort and might have little or no connection between them.

I wouldn’t rule it out as possibility, but I’d look very carefully at the two texts before I’d make anything of this. There is a tendency for there to be many claimants about the origins of great cultural documents such as Dante’s Comedy. I’d take the general claim of you Islamic professor with a grain of salt, and look carefully at what Dante scholars have concluded from their studies.

1On Wed, 30 Oct 1996, Darci Meline wrote: “In my Islamic class, the professor stated that in Dante’s descriptions of heaven and hell, he could have very likely gotten it from Mohammad’s account of a journey to heaven he was taken on where he was allowed to see the seven parts of heaven and the seven parts of hell. The professor said that this literature would have been translated into Dante’s language and that it was available to him. What do you think about that?”

6.2 Chaucer, Miller’s Tale

It’s not altogether clear what the term “clerk” means in this context^{2}, perhaps. The Middle English word develops in two directions: one evolves into “cleric” and the other into “clerk” (as in “sales clerk”). The first clearly has official, religious connotations; the latter is more of a job-position.

Both Absalom and Nicholas are clerks, but the question is whether either of them is an ordained “cleric” (with expectations of virtuous, celibate life). Clearly, neither is a priest/clergyman, though both may be at lower levels of status/education which might eventually lead to their being ordained (receiving the sacrament of Holy Orders). To that extent, then, they are still “lay” people and able to act accordingly.

Presumably, Absolom as a parish clerk would have been expected to behave himself better than some others might, but there doesn’t seem to be anything he does which goes over the edge much. In fact, many of Chaucer’s ordained clergymen behave even more “loosely” than Absalom. So if he is a venial sinner, the others are well on their way to “mortal” sin.

Anyway, how many of us really manage to live up to the ideals we (and others) profess we ought to live by. The Miller’s Tale is a comic, and somewhat exaggerated, somewhat cartoon-like work. Maybe these “serious” questions are a bit off the mark?

I have some difficulty getting a clear sense of what you are about here, but the essential question about how one becomes a saint is clear–even if it is not easy to answer.

In the early church, individuals who developed a reputation for holiness or who were killed because they professed the Christian faith were believed to be in heaven and therefore were honored by those who lived after them. This continued to be the pattern by which people “became” saints for much of the Middle Ages: a reputation for holiness, achieved by some public witness or actions, would result (over time) in an individual’s fame spreading beyond the local communinity where s/he lived, and achieving some more widespread recognition as a saint.

2On Sat, 2 Nov 1996, R. Disario wrote: “I just finished about 2/3 of the Miller’s Tale, and the character of Absalom is confusing me in one respect: how could he be a parish clerk and basically get away with the behavior that he does? Maybe I misunderstood the idea of a parish clerk, but it seems that his carrying on’s (the barhopping, the schmoozing of other men’s wives) would have been contradictory to the church ideals. Or, were the rules at this time different? I mean, he was committing the mortal sin of coveting the carpenter’s wife (so was Nicholas, but that’s a different question). I guess I’m mostly curious of how the church would’ve reacted. Chaucer doesn’t expand on that at all.”

As the Church became more centralized, however, it became more difficult for an individual’s local reputation to determine their wider recognition as a saint. As a result, greater “signs of sanctity” would be required: miracles, great learning, evidence of particularly unusual holiness, etc.

Lots of early bishops and teachers of the Church became saints of the wider church because, like Augustine, they effectively argued against heretics and enemies and defined the orthodoxies of the faith, which the Church as a whole adopted. These “confessors” of the faith (holy men and women who didn’t have to die for their beliefs) make up a large part of the “collection” of saints in the Middle Ages. But doing some particularly “huge and amazing” thing is not really a requirement–or perhaps I should say that the definition of “huge and amazing” is subject to change and adjustment in various circumstances.

6.3 Song of Songs

I apologize for our shorthand^{3} on “filioque.” I should have introduced it a bit more fully. The issue is one related to Trinitarian controversies: i.e., how the Father,Son and Holy Spirit are related to one another. It is usually fairly clear that the Son “derives” (or “proceeds” from the Father, but where does the Spirit come form. Some held that the Spirit proceeded from the Father alone (as the Son did) and others argued that the Spirit proceeded from both the Father and Son. So, in fixing the language of the statement of the Church’s essential beliefs (in a “creed”) the presence of “filioque” was a crucial point of disagreement in the passage declaring the “origin” of the Holy Spirit. Some would hold out for “from the father” while others would insist on “from the father and the son” as the language of the creed. On that point, there was a major split that virtually divided the church into two parts (East and Wes), so the reference to “filioque” is a shorthand way of referring to the fundamental effects (apparently) minor disagreements over language (and ideas) have had in the history of Christianity (and western culture).

3On Mon, 4 Nov 1996, C. Christian wrote: In class today towards the end Gabrielle jumped in and asked you if you had already discussed Filio-que. When you said no she went ahead and gave a quick description of it. All that I really understood was that it literally meant “from the Son-and.” The rest of the description I either didn’t hear or didn’t understand. I was hoping that you would be able to e-mail me back and give me a more understandable explaination. Also I didn’t understand where this idea came into our discussion. Please help clear this up for me.

Also I was hoping that you could give me your opinion of what you think the sexual/erotic ”Song of Songs’ is doing in the Bible. I know that you wanted us to think about it for ourselves, but I never really heard your opinion and greatly desire to.

I trust my comments in today’s class have revealed at least some of my view of the what SS is doing in the Bible. I’m not sure I have a simple direct answer (for this or for anything else), but I think the basic reason for its inclusion is that it was early on attributed to Solomon, and his reputation ensured its preservation and authority. How exactly the “meaning” of the Song might have fit into early conceptions of what the canon taught is not clear to me, but there is good reason to think that the marriage imagery elsewhere in the OT might well have allowed the erotic potential of the Song to serve more abstract cultural or theological purposes.

6.4 Rooster Poem

Modernized spelling:

I have a gentle cock, Crows me day; He does me rise early, My matins to say. I have a gentle cock, Comes he is of great; His comb is of red coral, His tail is of jet. I have a gentle cock, Comes he is of kind; His comb is of red coral, His tail is of ind. His legs be of azure, So gentil and small; His spurs are of silver white, Into the wortewale. His eyes are of cristal, Locked all in amber; And every night he perches him In my lady’s chamber.

Translation:

I have a noble^{4} cock, Who crows for me at daybreak; He causes me to rise up early, My morning prayers to say. I have a gentle cock, Who comes from a great family; His comb is red coral, His tail is jet-black. I have a gentle cock, Who comes from a noble family; His comb is red coral, His tail is indigo. His legs are azure, So fine and shapely; His spurs are silver white, Down to the roots. His eyes are cristal, Set entirely in amber; And every night he perches himself In my lady’s chamber.

6.5 Dante, Vita Nuova

Dante’s view^{5}, like that of the European Middle Ages generally, is that human nature is basically good, but fallen. Desire (love) is expressed as the attempt of that good nature to overcome the effects of the fall. That is clearly consistent with Boethius’s views. Pleasure is connected to larger concepts of “happiness,” but (as Dante seems to imply in the second section of Vita Nuova, and states quite explicitly in Canto 17 of Purgatorio) we must be ready and willing to make distinctions between (or among) the various kinds of pleasure/love/ happiness: “vital,” “animal,” “natural,” and “rational.” If you limit “pleasure” or “love” to the realm of the first three, then you are excluding what for Dante and other medievals (and still others at other times) would be the most ”real” level at which pleasure/love becomes human. If humans are rational animals, the Middle Ages, as inheritors of Graeco-Roman thought, will place the emphasis on the rational, on reason. If you extend the ”pleasure” that drives us as humans to that level, then your statement below would be, I believe, perfectly consistent with Dante’s.

4?yuppie 5In Canto 17 in Purgatorio, Dante writes, “The soul at birth, created quick to love, will move toward anything that pleases it, as soon as pleasure causes it to move.”

6.6 31 rules from the Art of Courtly Love

The Rules of Love from Andreas Capellanus (The Chaplain, De Arte Honeste Amandi) The Art of Courtly Love, Book Two

1. Marriage is no real excuse for not loving.

2. He who is not jealous cannot love.

3. No one can be bound by a double love.

4. It is well known that love is always increasing or decreasing.

5. That which a lover takes against the will of his beloved has no pleasure.

6. Boys do not love until they arrive at the age of maturity.

7. When one lover dies, a widowhood of two years is required of the survivor

8. No one should be deprived of love without the very best of reasons.

9. No one can love unless he is impelled by the persuasion of love.

10. Love is always a stranger in the home of avarice.

11. It is not proper to love any woman whom one should be ashamed to seek to marry.

12. A true lover does not desire to embrace in love anyone except his beloved.

13. When love is made public it rarely endures.

14. The easy attainment of love makes it of little value; difficulty of attainment makes it prized.

15. Every lover regularly turns pale in the presence of his beloved.

16. When a lover suddenly catches sight of his beloved his heart beats fast.

17. A new love puts an old one to flight.

18. Good character alone makes any man worthy of love.

19. If love diminishes, it quickly fails and rarely revives.

20. A man in love is always apprehensive.

21. Real jealousy always increases the feeling of love 22. Jealousy, and therefore love, are increased when one suspects

his beloved

23. On who is afflicted by love, eats and sleeps very little.

24. Every act of a lover ends in the thought of his beloved.

25. A true lover considers nothing good except what he thinks will please his beloved.

26. Love can deny nothing to love.

27. A lover can never have enough of the solaces of his beloved.

28. A slight presumption causes a lover to suspect his beloved.

29. A man who is afflicted by too much passion usually does not love.

30. A true lover is constantly and without intermission possessed by the thought of his beloved

31. Nothing forbids one woman being loved by two men or one man by two women.

(From the translation of John Jay Parry, edited by Frederick W. Locke, and revised by Mcel F. Vaughan)

Scholars and readers are really divided over whether Andreas’s ”rules” are serious^{6}. Some think they are; others think they are presented with tongue planted firmly in his cheek. I suspect they are ”serious,” but in a descriptive rather than prescriptive way: ”rules” derived from readings and discussion of ”love” in the courtly (and ecclesiastical) community Andreas moved in. (He was a chaplain to the Countess Marie of Champagne. And she was a very literary, cultured woman.) I suspect there is some humor in the rules (and in the book he puts them in), but I’m not sure we moderns will necessarily ”get” all the jokes he may have in there for his contemporary audience.

If you as so much in love with someone that you want them to ”be yours,” to want to have them all to yourself, wouldn’t that be a reason to feel jealous when you saw signs that that might not be the way it is. I agree that trust between people is an important part of a love relationship, but ”relationship” is not really what Dante and Andreas are dealing with: they are focused on the refinement of (male, primarily) desire (or obsession). Love is, for them, something that can exist in its perhaps purest form in a desire that never expresses itself in sexual fulfilment–or even in real reciprocation of that ”feeling” of love. The perception of goodness, beauty, etc. can inspire love. And it’s that, more than ”possession” that seems at the root of this so-called fin’ amors, or ”courtly love.” There are, no doubt, sexual elements to it, but for Dante those elements are almost refined away from the pure essence he celebrates.

6On Mon, 11 Nov 1996, J. Peper wrote: My questions pertain to the ”The Rules of Love” that we discussed last week. First off, it seems to me that many of the rules are more of a personal opinion than a justified (meaning everyone would agree) rule. I realize that not everyone is going to agree with anything that is set in front of them. Now, some examples of what i mean by a personal opinion is that there are several messages that deal with jealousy and love. In my own personal opinion, I believe that jealousy should be put away when a person encounters love. When a personal falls truly in love with another individual shouldn’t he or she trust? Isn’t trust a huge part of love? It was never mentioned in the 31 rules. Another question that i have is with rule number 11. When people are young and they fall in love it seems like it would be hard to ”want” to marry.

That too is just a personal opinion. Overall, I’m asking if this love rule list is really real?

Do some people live by every rule on here?

I’m not sure your response to rule 11 puts the emphasis in the right place. I take that to mean that one should only love a partner whose class and reputation is equivalent, or superior, to your own. Being not ashamed to marry is not the same as wanting to marry.

6.7 Chaucer’s Audience

Chaucer’s immediate audience was his friends at the court of the English king, Richard II (last two decades of the 1300s) and in the ”civil-service” of London and Westminster of that time. It’s not clear how ”The Miller’s Tale” was received by the church, but the manuscripts of the Canterbury Tales are rather numerous, and some of them, no doubt, were copied by clerics–though not all, since there were ”professional” scribes outside of monasteries and cathedrals at this time. But the Tales were for a most secular, merchant- and upper-class audience–which is not to say they were not ”good” Christians also.

I would, by the way, make a sustained argument that in many important respects, this tale of the miller is one of Chaucer’s most ”moral” tales–despite (and perhaps even because of) its bawdiness and physicality. Just as the ”style” of the ”Roland” and Einhard’s ”Life” can provide no sure guide to their ”accuracy” or truth, so the apparent irreverence of the ”Miller’s Tale” should not distract us from its deeper meaning. That tension between surface and depth, literal and allegorical is at the core of the problems I was having us investigate this week in regard to the ”holiness” of the “Song of Songs.” Does Dante’s “Vita Nuova” present, do you think, a truer image of human love than Chaucer’s tale does?

6.8 Miller’s Tale

These ideas^{7} are, as you’ll see, the opinions of working-class men (the Miller and Alison’s husband John), who both to be very happy with such wilful ignorance as a way of maintaining his own positive self-image in the face of Nicolas’s pointy-headed intellectualism and the Reeve’s moralism. There is a blue-collar anti-intellectualism here that we see in our own time, and I think that is Chaucer’s point here: to establish clear class (and individual) distinctions among the people on his pilgrimage to Canterbury and in the “Miller’s Tale.” In one sense, this tale is a response to the “romance” of the preceding “Knight’s Tale” in which the two cousins pursuing Emily are almost indistinguishable as lovers or people. That’s certainly NOT the case in the Miller’s version of a very similar story.

There may be self-interest in certain quarters of societies (then and now) involved in keeping certain classes (or genders) undereducated. But it is also the case that some people have a much higher opinion of their knowledge and wisdom, their understanding of the world and how it works, and the superiority of their “morality” than others would accept. The “wisdom” in these two quotations would, I expect, not have been endorsed by Chaucer– except as accurate representations of the comic limitations of some people and their self-satisfied confidence in themselves and their opinions.

I expect that Chaucer, though not a monk, would have been much more on the side of the views you attribute to monks here: he was an avid student of classical and contemporary writings, he was a politician and diplomat of some standing, he was an accomplished government administrator, and he was very interested in the science of his day, particularly in astronomy (he wrote a book on the Astrolabe, and perhaps also another one on planetary cosmology). So he clearly would NOT have sided with John’s condemnation of scholars who spend their time inquiring into “astromy.”

6.9 The Summa

The “Summa^{8}” usually refers to the huge compendium of Thomas Aquinas^{9}, a thirteenth-century theologian at Paris who examined almost the complete range of theological doctrine in an analytical and systematic fashion. He is usually represented as summarizing most of the important developments in theology and in logic that flourished in the late=twelfth and early-thirteenth centuries under the influence of Aristotelian ideas and methods. While some of his positions were not at first considered orthodox and he had to recant some of his views, by the time of the Protestant Reformation of the sixteenth century, Thomas’s Summa Theologica (or Theologiae) had achieved a position very close to being THE orthodox statement of Roman Catholic theology.

7Re: Sarah Millers comment

8On Tue, 12 Nov 1996, Mckenzie Nielsen wrote: What is the difference between summa and Summa, I feel like I missed something, if there is a document known as the Summa, who wrote it, where did it come from?

9See also section 1.1 on page 1

(I’m not sure what distinction you are making between summa and Summa, but I suspect it is likely to be one that distinguishes between Aquinas’s Summa and the category of encyclopedic textbooks that are often called by that name. A summa, then, would be one of those works, while the Summa would refer to on particular text, such as Aquinas’s.)

Your sense of the cultural shift from epic to romance strikes me as being very astute. There is something quite significant occurring during the twelfth century (or thereabouts) that represents a more settled and “civilized” world, one less dependent on military power (at least within the “settled” realms of Europe). What military activity is encouraged is abroad, on crusades for example. There is some real effort to resolve differences among European lords, etc., is more judicial and negotiated ways. Though of course there is still fighting going on, it is more small-scale and local, I think–at least until the “national” wars of the fourteenth and fifteenth centuries start up.

The invention of “romantic” love in Europe during the 12th century is a recurrent thesis. My view is that this is mostly true, though of course there were practices that look very much like it in earlier and other cultures. The “Song of Songs,” for example strikes me as having many of the civilized and civilizing hallmarks of “courtly” love, and there are continuing arguments about whether it was really “invented” in 12th century Europe. Some point to traditions in Islamic poetry, for example, and even to things in ancient Greek and Roman poetry, which could be seen as “origins” for these ideas. IN any event, something new does seem to be rising up in the twelfth century in So. France (and spreading from there), whose influence continues. Whether this is only the result of the poems being now written down (and copied later) for the first time, or whether these sorts of poems are only being composed for the first time will keep scholars busy for another century or so.

One recent thesis, put forward very convincingly in a couple of books by Stephen Jaeger of our UW Germanics Department, which I mentioned in class, is that the language of this poetry is strongly influenced by and even derived from the highly rhetorical language of male “friendship” in schools and other courts during the late Carolingian period. You might want to take a look at Jaeger’s books if you’d like to pursue these ideas.

The bibliography of “courtly love” is a rich and varied one, full of interesting forays into various corners of literary and social culture in the Middle Ages, but also one where contesting (and sometimes quite acrimonious) views battle for ascendancy. It can be a “dangerous” place, so if you’d like any guidance from me (who of course has NO prejudices or biases), let me know.

Types of Women - Eva Ave

7.1 Ruth

The Old Testament^{1} for the most part does NOT hold that all people were the same. Again and again, discriminations are made between the good (or chosen) people favored by Yahweh, and those sent off into the wilderness (or elsewhere). Cain and Abel, Isaac and Ismael, Israelites and Canaanites, etc.

One of the oddities about ”Ruth” is that it has her come from the ”despised” tribe of the Moabites. There is some real intent here to be a bit more broad-minded, and by introducing Ruth into the line of David’s ancestors, the writer of this book clearly wants to take a more universalist, a less ”we’re the pure, chosen people” line. (Aren’t there some today, I believe, who still hold that you are only a Jew if your mother is Jewish.)

The identification of blood lines with ”virtue” is a long-standing, and repeatedly countered view, and as Gabrielle said in Wed. class, it is an element in the Wife of Bath’s Prologue also. And clearly Jesus (in the gospels) and Paul were not bound by such ethnic or racial views of the Church (the new chosen people). Jesus’s interaction with Samaritans (incl. the Wife of Bath’s analogue, the woman at the well) strongly insists on God’s being a universal God. And the debate between Paul and Peter on the treatment of Gentile converts (in Acts) marks a crucial stage in the growth and direction of the early Christian Church.

1On Wed, 13 Nov 1996, S. Cantwell wrote: In class yesterday we were talking about how non-Israelite women were portrayed. Delilah (sp?) was a bad example and some were quite good. I was wondering why, if God/Jesus thought that all people were suppossedly the same, that this segregation existed in the text. I guess it can possibly be explained by translations, but it just seems strange. It is also possible that I am looking at the example from the 20th century version of religion.

7.2 Judith

7.2.1 Judith as a Woman

This^{2} is clearly something of a ”problem,” I agree. It makes me think that the author of ”Judith” has a strong sense of ends justifying means, and even apparently immoral or unsavory actions, when taken in the service of a good cause, are justified and not subject to condemnation for not complying with the strictures handed to Moses on Sinai.

The Anglo-Saxon poem clearly represents Judith as weaker physically than a man, by having her have to hack at Holofernes’s neck a couple of times before she can get the head off. That’s not in the Bible account. So divine grace or inspiration is clearly on the mind of this poet, and he gives it some prominence.

7.2.2 Murder as Justified

David’s comments rightly raise the issue of when ”murder” (or ”killing”–I’m not sure I make the same distinction as David does between then) may be ”justified.” We have no problem, in our own legal system, in find certain murders ”justifiable homicide”–as for example in cases of self-defense. The idea that some killing is ”justifiable” clearly underlies much of the Old Testament, and is by no means limited to ”Judith.” Think of all the ”holy” wars the ancient Hebrews and Israelites were involved in, killing enemies, exterminating tribes that were in their way to the Promised Land. There must have been some heavy footnotes explaining the nuances of ”thou shalt not kill” that allowed such behavior. In other words, a literal, absolute interpretation of that commandment is not by any means maintained in the Old Testament.

I didn’t read Justin’s earlier posting as sarcastic, but that may be my notorious problem with ”tone.” Much of what he says, however, whether he agrees with it or not, is quite compatible with ancient (e.g., OT) and medieval (e.g., crusades) views. Some killing is clearly permitted–and even encouraged–despite the fifth commandment. Some can be justified on individual grounds, and we seem willing today to accept those. Others are justified on social or ideological grounds, and we may be less willing to accept those. But aren’t ideas of ”national security” communal equivalents of ”self-defense,” extending the ”self” from the individual to the social? Few wars the US has engaged in have been justified on the grounds that the rule against killing has been retracted.

2Then in the fight had Judith won herself outstanding glory, as God granted her when heaven’s prince gave her the victory.

In much of the Bible and of medieval European culture there were reasons which permitted killing.

Am I wrong in thinking that our focusing on this in the case of Judith’s killing of Holofernes reveals some particular concern with HER actions, especially her blatant acceptance of the need for killing as the solution to her town’s difficulties, and her confidence that God would not only endorse her act, but provide the assistance necessary to carry it out?

7.3 Numerology, astrology, non-Biblical texts

It’s not so much that they^{3} are treated as ”divine science,” but that understanding creation is one sure way of understanding the Creator. It’s only when numbers or the stars/planets are perceived as opposed to what is accepted as ”divine science” (or theological truth) that problems arise–as in the case of Galileo, or the recent debates about ”evolution.”

If one believes firmly, as most medievals did, that the cosmos was ordered by God, then understanding that order (whatever it turns out to be) is an effort worth undertaking. If, as in our own more modern times, there is a disjunction (if not downright opposition) between ”science” and ”theology,” then something radically new has occurred than was known in the Middle Ages. As we’ve seen (I hope) in this class, the opposition of a LITERALLY understood Bible and a SCIENTIFICALLY understood cosmos is really not evident in the M.A. Indeed, even the concept that there could be such an opposition is alien to the period. As I have repeatedly said (and as we saw early on in Augustine’s ”On Christian Doctrine”), a simple, LITERAL reading of the Bible is NOT what one comes to expect (or value) in medieval texts. The allegorized, spiritual understanding of the text is favored, in part because only at that level can the patent inconsistencies among the various texts contained in the Bible be resolved..

3On Sun, 17 Nov 1996, Sarah Fisher wrote: in reading the biblical excerpts for the week, matt 1 talks about fourteen generations between abraham and david, david and the exile, and the exile and christ... other things we have read concern themselves with numerology. do people of the middle ages believe numerology to be a divine science? as well as astrology, since the wife of bath discusses the importance of both mars and venus?

where is the division between godly beliefs and heresy?

In part because of the focus^{4} of this course’s topic and themes, I’ve not had you read many things that are not in SOME way connected (or at least connectible) to Biblical issues. There are, of course, many things that are not directly (and sometimes not even indirectly) connected to the Bible. Some of stuff we’ve read (e.g., Boethius, the Old English poem ”Deor,” and even the ”inner” sections of ”The Wanderer” -leaving off the opening and closing lines-) are only conceptually related to issues we discussed from a Biblical perspective.

There is a lot of stuff that is not ”Biblical” in any overt sense, though of course they were likely copied and preserved by monks and other religious, and would probably not have been so preserved had they been seen as opposed to the kinds of values that Christianity upheld. And since most of THOSE have some connection to Biblical texts, everything that was copied by monastic scribes probably could be seen as in some way connected to those texts. (And similar claims might be made for the rich textual traditions of Jewish writings, which we have not looked much at, in the Middle Ages: they also issue, for the most part, from religious environments also.)

And all of this leaves aside the body of legal and political and ”historical” material that provides records of events from the Middle Ages. Some, but certainly not all, of these bureaucratic texts bear little if any overt connection to the Bible.

7.4 Second Shepard’s Play

This^{5} is a “class” complaint: that is, the shepherd is complaining about how the ”upper classes” treat him and his lower-class colleagues. There is something of a democratic spirit to these shepherds. As you see, when Mak enters, he tries to pretend his is ”above” these shepherds, that he is the messenger of a great lord. The theme is not a major one in the play, but since it is a play about how God humbled himself to become human, it may be a not-altogether-incidental element. In the so-called ”Peasants’ Revolt” of the 1380s in and around London, there was a saying that seems to capture some of this sort of anti-classist thinking: ”When Adam delved and Eve span, Who was then the gentleman?” In other words, since humans were created equal, why do some think they are better than their fellows. This ”Scriptural” democratic spirit works against clergy and the secular class-structure, and late-fourteenth- and fifteenth-century England was a hotbed of such thinking. It expresses itself somewhat later in the Puritan Revolution in England–and in the earlier emigration of Puritans (and other dissenters) from England to the American colonies. So, as you see, this stuff we’re reading is DIRECTLY connected to our modern circumstances.

4ibid. as well, you said that The Second Shepherds’ Play was part of a biblical cycle of plays... and all of the works we’ve read thus far have specific biblical and religious purposes. was there any small amount of works in the middle ages which didn’t apply to the bible in any manner? do we just not have any record of them because they weren’t written down by the monks? or did they not exist?

5On Sun, 17 Nov 1996, T. Becker wrote: In the Second Shepard’s Play at the beginning it says: “We are hand-tamed By these gentle-men.” I was wondering who are the gentlemen that are being refered to?

7.4.1 The role of Mac

Mak functions^{6} as something of a parodic ”angel,” perhaps even as a sort of humanized devil-figure. But he is also related to the shepherds in some ways, and how he is treated at the end of the play is something we will want to discuss in the Thurs/Fri sessions.

If you look at the way the play is structured, you can see that there are two specific houses set up: one for Gyll, Mak, and their ”child” and another for Mary, the Angel, and ”their” child. The action of the play moves from the site of a false ”nativity” to that of the true Lamb of God. The play makes rich use of the ”literalizing” vs ”spiritualizing” of sheep and shepherds.

7.5 Marriage

By this point^{7} (early 15th century) the Church would not even condone, much less permit, any divorce. In fact, as I said in class earlier in the quarter, that is still the position in the Roman Catholic Church. Before the 10th or 11th century (or thereabouts) marriage was not really a ”church” event, and it was not really treated as a ”sacrament.” The legal/economic relationship was subject to secular or civil or common laws in the various cultures of early medieval Europe, and it is only afterwards that there evolves ecclesiastical sanction.

6On Sun, 17 Nov 1996, C. McVicker wrote: I ”The Towneley Second Shephards’ Play”

what purpose did the portion containing Mak serve in the entire play. It seemed to be seperate from the angel informing the shephards of Jesus’ birth. However I thought that it must have had a purpose in the eyes of the author. Is there something I am missing that ties the two together?

7On Tue, 19 Nov 1996, J. Reade wrote: In ”The Book of Margery Kemp”, Margery is having major maritral troubles with her husband because she wants to become chaste, and he insists on his rights. In situations such as this during this period, was the church willing to grant divorces to wives under any circumstances?

Chaucer’s Wife of Bath, after all, talks about weddings ”at the church door”: even as late as 1400 marriages were still not conducted IN the church. And to this day, the theology of marriage in the Roman Catholic tradition is that the ministers of marriage are the husband and wife, not the clergyman/priest, who merely ”witnesses” the exchanged of vows on behalf of the church. It is, in other words, still conceived of as a contract between man and woman, and it is only because they are Christian that this contract has sacramental power.

Birth and Redemption - Week 8

Chaucer, Prioress’s Tale and Clerk’s Tale; “Second Shepherds’ Play”, “Abraham and Isaac”

Isaiah 6:1 to 8:4; 11:1-16; Matt 1:1 to 4:25; Luke 1:1 to 2:52; John 1:1-34

8.1 The Book of Margery Kempe

At least one of the reasons^{1}, aside from her ”personality,” is that there are some elements of ”anticlericalism” in her ”mystical” relations with God. Those who have (or claim) direct conversations with God, unmediated by ecclesiastical structures and clerical surrogates, pose a direct criticism (attack?) of the established institutions.

The accusation repeatedly leveled at Margery, that she is a ”Lollard”– even if it is always finally rejected–is one that was important in the early decades of the fifteenth century as a source of ecclesiastical and even secular danger in England. There are numerous heresy trials and recantations and heretic-burnings of Lollards, who criticized secular collusion with ecclesiastical errors. The revolt of Sir John Oldcastle (one of the models for Shakespeare’s John Falstaff in the Henry plays) resulted in his execution.

1On Mon, 18 Nov 1996, Libby Charhon wrote: I’m really curious as to why some of the clergy had such an aversion to Margery Kempe. I know that she probably wasn’t the most pleasant person to be around, but why label her a heretic, etc.? Would they (the clergy) have treated a man professing to have the ability to speak with christ in the same manner? Were bishops and other church leaders so paranoid in the Middle ages? Or did they just feel threatened by a woman claiming to have ”a gift” that they themselves did not possess?

While women were prominent in the Lollard movement, they were hardly singled out for harassment. Indeed, it might be said that men were more suppressed for their involvement. It was a movement with roots at Oxford University (among clerics, and John Wyclif is a famous name associated with some of their ideas), so it was not a marginal movement. It has some connections with the ideas and beliefs of dissenters in the English Renaissance, like the Puritans.

Margery in many ways and in many of her opinions sounds very like Lollards, but she never goes as far as they do in rejecting existing secular and religious institutions and authority. She may criticize and argue with them, but she is not dismissive. Her unmediated contact with God comes close to being an orthodox mystic’s religion. On the surface it has elements that closely correlate to those of heretics, but the fundamental point of difference is that heretics reject the institution and its authority; Margery doesn’t. Her anticlericalism is mild, and mostly leads to reformative criticism. That some of the clerics don’t see that as a positive ought not to be surprising. We don’t like to have our faults and failures pointed out (loudly) in public.

8.2 The Prioress’s Tale

8.2.1 Anti-Semitism - G. Shearer

I thought I’d reply to this en masse, since some of the confusion^{2} re The Prioress’s Tale might be cleared up this way. Part of the Tale, which hopefully will clarify the boy’s death, was omitted accidentally. This was my fault, and I have corrected it. The story is now intact, and should make more sense. The story’s anti-semitism cannot be edited or amended, however, and should be addressed. One thing to keep in mind, as Miceal has suggested, is the character of the person telling the tale. The Prioress is a shallow, pretentious person, who enjoys the status gained through assuming holy orders. It seems in character for her to tell a maudlin, and what is to us highly offensive, story.

2On Fri, 22 Nov 1996, D. Thomas wrote: On Friday we did not get to any discussion of the Prioress’s Tale, so I would like to ask a late MMQ about it. No one has yet said anything about why there is no explanation of the murder of the child. It apparently takes place between the 7th and 8th stanzas. Why does Chaucer not go into detail about why the boy was killed, or who exactly killed him? I know by the strong anti-semitism that it is supposed to be a Jew, and the murder is followed by a small version of the Holocaust, but is the murderer supposed to be just any Jew? Don’t individuals count at all? I’m reading this as simply a Jew-bashing anti-semite story, and am still wondering about the circumstances of the murder itself.

The middle ages as a whole was not known for religious tolerance, however, and it might be too much of a stretch to say that the story would be offensive to the medieval reader in the same way that it is to us.

8.2.2 The pearl - Vaughan

Your characterization^{3} of the tale’s view of Jews is pretty accurate, but you will note that it is set off in Asia, not in Europe, so it may have no direct relevance to the historical realities of Jews in Europe. The tale’s teller is a cloistered Englishwoman (of somewhat limited intelligence and sympathy, I would say). And the author, Chaucer, while he may have had contact with Jews in his travels (or in his work as a customs official in London), was living in a country where there were “officially” no Jews for about a 100 years.

The lord of the tale retained the Jewish ghetto for, as the translation has it, “their foul lucre, by usury gained.” In other words, they are there because of their money and their “banking” services. (There were, for most of the Middle Ages, specific rules against Christians’ lending money at interest, and so various “foreigners” functioned as bankers and moneylenders as the economy became more and more based on monetary exchanges.)

About the “pearl”^{4} that keeps the boy alive: the poem (it’s not a play) doesn’t make an explicit connection with the communion bread of the Mass, but you are right, I think, to see a possible connection. (I always wonder, when I read this tale, whether we are supposed to be critical of the abbot who takes it from the boy’s tongue, which results in his actually dying. Why didn’t he leave it there so the boy could continue to live and sing ”Alma Redemptoris Mater”?)

3On Mon, 18 Nov 1996, Victor Cuanzon wrote: From reading “The Prioress’ Tale,” I got the impression that the relationship between Jews and Christians during that time (or at least that town) was not good. The Jews of the town were tolerated, but I don’t think they were accepted. My question comes from the beginning of the tale where it seems to say that the one quarter Jewish population was preferred by the lord (kept there by force?). What purpose did they serve for the lord?

4Ibid. Another question I have from the same tale concerns the end of the play where the boy stays alive and singing as long as he has the grain on his tongue. Is the grain supposed to represent bread –i.e. the Body of Christ? Or is the grain a symbol of something else? Or nothing at all?

Death and Salvation - Week 9

Chaucer, Pardoner’s Tale

Matt 26:1-end; Mark 14:1-end; Luke 22:1-end; John 18:1-end; Job; Isaiah 36:1 to 55:13;

9.1 Chaucer Pardoner’s Tale

9.1.1 The post-story requesting of alms

Requesting ”alms” (to support his ”hospital”–and himself) is clearly one of the purposes of preaching indulgences of the sort he offers. While they were, strictly speaking, not ”sold,” money-raising for worthy causes were what these indulgences were authorized for. Of course, there were unscrupulous and dishonest ”pardoners,” and Chaucer^{1}, clearly raises doubts in our mind about this particular pardoner’s morality and honesty.

His tale^{2} is told to inspire in hearers the need to seek grace to protect them from the punishments of Purgatory after death. And he has just the ticket, in the ”plenary” indulgence he offers. The fact that he muddles the distinction between guilt and penalty (”culpa” and ”poena”) may complicate matters more, but that is something for us to discuss more in class tomorrow.

1See section 6.7 on page 24 for a discussion of Chaucer’s Audience

2On Sat, 30 Nov 1996, Suzanne Renner wrote: I was wondering why the Pardoner asked for money to pardon people after he told the tale. Is that to show he’s not as pious as he professes? Or he tells the story to encourage the people to feel guilty?

9.1.2 Relationship between pardoner and priest

A pardoner is not a priest, and is not an official member of the clergy at all. (Some may have had church training/education, but I am not aware of any priests functioning as pardoners.) He is a sort of church-approved fund-raiser employed by various church bodies (hospitals, shrines, churches, etc.). Officially, he is authorized to distribute an indulgence and to request alms (donations) from the people he talks to. Since there is the (relatively easy) possibility of this being presented or seen as a commercial transaction– exchanging money for ”pardon,” earthly goods for spiritual ones–it was in fact a source of abuse and of major criticism of the Church in the late Middle Ages. (One might even say that it was the lightning rod issue for the Protestant reformers, and one which was a not-very-easy-to-defend-against criticism.)

Your larger questions^{3} are not as easily answered, of course, and I am hesitant to generalize about a huge and complex intstitution on the basis of (even widespread) individual ”sins.”

3C. McVicker wrote: In the intro to the Pardoner’s Tale the pardoner, which I assumed was another name for a priest, talks of being in it for the money. He freely says that he is corrupt and uses fake relics and doesn’t follow his own preaching etc. I initially wondered if this type of deception was a common problem during the time period. But then I got to thinking and wondered would the corruption of a priest have a negative impact of the people around him, because he would still be doing the same thing only with a different mindset on the issue. I guess what I am trying to ask is how big of a problem was corruption during the time period and what impacts does it have on the community around it. The reason I am interested in this is because in the earlier chaucer stories he made a point about the speaker through the speakers outlook on life.

3http://weber.u.washington.edu/d52/cl280

The Otherworld and Afterlife

Apocalypse and Revelation

Revelation; II Cor 12:1-11; Ezech 1:1 to 3:21; and selections from I and II Thess; Daniel; and Zechariah